The triangles have the same perimeter. the first triangles measurements are X+3,X+3,X+3
The second triangle X+3,X,2X-3
What is the value of X?

Answer by cleomenius(959)

x + 3 + x + 3 + x + 3 = x + 3 + x + 2x -3
3x + 9 = 4x
9 = x
12 + 12 + 12 = 12 + 9 + 15
36 = 36
Cleomenius.